### v1

# Structured Probabilistic Coding

Abstract
This paper presents a new supervised representation learning framework, namely Structured Probabilistic Coding (SPC), to learn compact and informative representations from input related to the target task. SPC is an encoder-only probabilistic coding technology with a structured regularization from the target label space. By extracting compact and informative representations from input related to the target task, SPC can enhance the generalization ability of pre-trained language models for better language understanding. Specifically, the hidden representation is encoded into a Gaussian distribution space, while maximizing the prior entropy of latent representations concerning label space. This technique can simultaneously perform information encoding and task prediction in one module to more fully utilize the effective information from input data, and use variational inference in the output space to reduce randomness and uncertainty. To better control the probability distribution in the latent space, a structured regularization is proposed to promote class-level uniformity in the latent space. With the regularization term, SPC can preserve the Gaussian distribution structure of latent code as well as better cover the hidden space with class uniformly. We conduct evaluations on 12 natural language understanding tasks. The results show that our SPC can effectively improve the performance of pre-trained language models for various classification and regression tasks. Experiments demonstrate that SPC can enhance the generalization capability, robustness to label noise, and clustering quality of output representations.
Introduction
Probabilistic embedding (Vilnis and McCallum 2015) is a flexible representation learning technology whose goal is to learn the underlying probability distribution of data. In contrast to deterministic embedding (Pereyra et al. 2017; Miyato, Dai, and Goodfellow 2017; Gunel et al. 2021), which maps each data to a fixed vector representation, probabilistic embedding embraces the notion of learning a probability distribution, mapping each data point to a distribution. These probabilistic embedding approaches can better describe the uncertainty and complexity of data, handle redundant information, and provide better discriminative representations. Probabilistic embedding has been applied to various domains such as computer vision (Oh et al. 2019; Shi and Jain 2019) and natural language processing (Mahabadi, Belinkov, and Henderson 2021; Hu et al. 2022a).
Most probabilistic embedding methods (Kingma and Welling 2014; Alemi et al. 2017; Higgins et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) are or can be built upon the information bottleneck (IB) principle (Tishby, Pereira, and Bialek 1999; Tishby and Zaslavsky 2015). The principle aims to find a maximally compressed representation of the input that maximally preserves information about the output, striking a balance between compression and prediction. These IB-based methods typically involve two parametric modules, i.e., an encoder and a decoder. The encoder maps the input data to a probabilistic distribution in the latent space, and the decoder maps the probabilistic distribution to class probabilities for classification, predicted output values for regression, and generated samples for generative tasks. The encoder and decoder are trained jointly to minimize a task-specific loss function and maximize the mutual information between the input data and the latent space.
However, under the encoder-decoder architecture, the process of mapping input data to probability distributions by the encoder may lose some task-related information which is essential for decoder. This is because probability distributions inherently contain randomness and uncertainty, which may be irrelevant to the task and interfere with the model’s learning of task-related information. To avoid this, we propose an encoder-only embedding technology, probabilistic coding, that combines probabilistic encoding and task prediction into one module. By using variational inference in the output space, we can better control and utilize randomness and uncertainty of data. This helps improve the model’s generalization performance, especially when facing incomplete or noisy labels. Combining probabilistic encoding and task prediction into one module can promote to learn more compact representations. These representations can better capture the underlying structure of data, thereby improving task prediction performance.
Besides, although probabilistic embedding methods can capture data uncertainty and complexity, they are restricted to limited or biased data, which cannot fully represent the true distribution of the target task. Therefore, in the process of mapping input data to probability distributions in the latent space by the encoder, some task-related important information may be missing to some extent. The insufficient information may lead to limited task performance on new data and poor model generalization. To improve task prediction ability of the latent representations, we leverage the structured information of target task space to constrain the learning process of the latent space’s probability distribution. Under the framework of probabilistic coding, the structured regularization of latent space can help the model learn more informative representations related to the target task, thereby improving the model’s prediction accuracy on new data.
In this paper, we present a new supervised representation learning framework, structured probabilistic coding (SPC), an encoder-only probabilistic coding technology with a structured regularization from the target label space. By extracting compact and informative representations from input related to the target task, SPC can enhance the generalization ability of pre-trained language models for better language understanding. Specifically, we use variational approximation to encode hidden representation vectors into Gaussian distribution spaces while maximizing the prior information entropy of latent representations concerning label space. This technique can simultaneously perform information encoding and task prediction in one module to more fully utilize the effective information of input data, and use variational inference in the output space to reduce randomness and uncertainty. The method employs probabilistic encoding technique combined with labels to fit the true distribution of data well and encode the semantic structure of label spaces. At the same time, introducing the structure information of target task space during the probabilistic coding process to constrain the latent space’s probability distribution can make the probability distribution of latent space better reflect task-related information, which helps improve task prediction performance. The structured regularization promotes class-level uniformity in the latent space under the multivariate Gaussian distribution. With the regularization term, SPC can maintain the Gaussian structure of the neighborhood in the input space while achieving the best possible coverage of the hidden space with uniformity across classes.
We conduct experiments on 12 natural language understanding tasks, including 10 classification tasks such as emoji prediction, social emotion detection, hate speech detection, irony detection, offensive language detection, sentiment analysis, and stance detection, as well as 2 regression tasks including semantic similarity prediction and plausible clarifications ranking. The results demonstrate that our SPC method can effectively improve the performance of pre-trained language models for classification and regression tasks. For instance, with the RoBERTa backbone, SPC can improve average performance by +4.0% and +1.5% for classification and regression tasks compared to CE. Our SPC framework consistently achieves the best average performance compared to other methods, including traditional deterministic methods (i.e., CE/MSE, CE/MSE+CP, CE+AT, and CE+SCL) and probabilistic methods (i.e. VIB, MINE-IB, and MEIB) under different backbone models such as BERT and RoBERTa. Extensive experiments show that SPC can enhance the pre-trained language models’ generalization capability including out-of-distribution and low-resource scenarios, robustness to label noise, and clustering quality of output representations.
The main contributions are as follows: 1) We propose a novel encoder-only probabilistic coding method that combines probabilistic encoding and task prediction into one module. 2) We design a new structured regularization term to promote class-level uniformity in the latent space for better task prediction ability of probabilistic embedding. 3) We propose a supervised representation learning framework, namely SPC, to learn compact and informative representations from input related to the target task. 4) Experiments on 12 benchmarks show that SPC gains state-of-the-art performance on classification and regression tasks. Extensive experiments demonstrate that SPC can enhance the pre-trained language models’ generalization capability, robustness to label noise, and clustering quality of output representations. [^1] The code is available at https://github.com/zerohd4869/SPC
Figure 1: Comparison of our SPC with existing deterministic embedding and probabilistic embedding methods.
Methodology
In this section, we present a supervised representation learning framework, structured probabilistic coding (SPC), to learn compact and informative representations from input related to the downstream task. As shown in Figure 1(c), SPC is an encoder-only probabilistic coding technology with a structured regularization from the target label space.
Probabilistic Coding
The probabilistic coding combines probabilistic encoding and task prediction into one module. Different from existing probabilistic embeddings applying encoder-decoder architecture, our encoder-only paradigm can effectively maintain task-related features and avoid negative effects caused by the randomness and uncertainty of probabilistic encoding.
Following Alemi et al. (2017), we employ a variational approximation to encode the input data into a Gaussian distribution space. Specifically, under the assumption that $p(t|x,y)=p(t|x)$, corresponding to the Markov chain $Y\rightarrow X\rightarrow T$, we aim to minimize the mutual information between the input $X$ and the latent representation $T$, i.e., $T\in\mathbb{R}^{|\mathcal{Y}|}$, as well as minimize the Kullback-Leibler (KL) divergence between the true posterior distribution $p(t|x)$ and the approximated distribution $r(t)$. The objective of probabilistic coding can be:
$\begin{split}\mathcal{L}_{PC}=\mathbb{E}_{t\sim p_{\theta}(t|x)}[-\log q(y|t)]+\beta KL(p_{\theta}(t|x);r(t)),\end{split}$
where $q(y|t)$ is a non-parametric operation, i.e., argmax function. $r(t)$ is an estimate of the prior $p(t)$ of $t$. $p_{\theta}(t|x)$ is the probability distribution for $x$, which is learned by the probabilistic encoder $\theta$. $\beta>0$ is a trade-off hyperparameter which controls the trade-off between the sufficiency of $t$ for predicting $y$, and the minimality of the representation $t$ from $x$. Here, $p_{\theta}(t|x)$ is typically set to the unit Gaussian $\mathcal{N}(0,I)$, i.e., $p_{\theta}(t|x)=\mathcal{N}(t|{\mu}(x),\Sigma(x))$, where $\mu$ and $\Sigma$ represent the mean and diagonal covariance, which are predicted by MLP with a single hidden layer. In probabilistic encoding, the re-parameterization trick (Kingma and Welling 2014) is used to make the model take gradients as usual.
In existing IB-based methods (Alemi et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) with an encoder-decoder architecture, their decoder can be a parametric approximation $q_{\phi}$ of $p(y|t)$. That is, the compressed representation $t$ can be sampled from the distribution $p_{\theta}(t|x)$, meaning that a specific pattern of noise is added to the input of $q_{\phi}(y|t)$. This noise would decrease the information conveyed by $t$ and enable them to lose some task-related information, which is essential for the learning of decoder $\phi$. Different from them, our probabilistic coding applies a non-parametric operation to predict, and combines probabilistic encoding and task prediction into one encoder module with the network $\theta$. It can effectively avoid negative effects caused by the randomness and uncertainty of probabilistic encoding.
Structured Regularization
As mentioned above, the Markov assumption restricts that the representation $T$ cannot depend directly on the target task $Y$. This means that the learning of $t$ does not fully utilize information of label space. Accordingly, the learned representation cannot sufficiently represent the true distribution of the target task, leading to poor generalization ability when learning from the limited or biased data. Therefore, we design a new structured regularization to explore underlying patterns of the label space.
Specifically, we add an additional term about the latent distribution to the objective function that maximizes the prior entropy of $T$ on the label space:
$\max H(T),\ \text{where}\ T\in\mathbb{R}^{|\mathcal{Y}|}.$ (2)
In the implementation, we utilize each sampled batch data to estimate $H(T)$. The Jensen’s inequality is first applied to obtain a lower bound, i.e.,
$\begin{split}H(T)&=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{|\mathcal{Y}|}p_{i,j}\log p_{i,j}\\ &\geq-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{|\mathcal{Y}|}\frac{1}{B}\sum_{k=1}^{B}p_{k,j}\log\left(\frac{1}{B}\sum_{l=1}^{B}p_{l,j}\right),\\ \end{split}$
where $B$ is the batch size. $|\mathcal{Y}|$ is the number of classes. $p_{i,j}$ represents the probability of sample $i$ belonging to class $j$ in the latent space $T$, which is computed by the output of $p_{\theta}$, i.e., the encoder of probabilistic embedding. Then, the Monte Carlo method is used to estimate this lower bound:
$\begin{split}H(T)\geq\hat{H}(T)&\approx-\sum_{j=1}^{|\mathcal{Y}|}\left(\frac{1}{B}\sum_{k=1}^{B}p_{k,j}\right)\log\left(\frac{1}{B}\sum_{l=1}^{B}p_{l,j}\right)\\ &\approx-\sum_{j=1}^{|\mathcal{Y}|}\overline{p_{\cdot,j}}\log\left(\overline{p_{\cdot,j}}\right)=\mathcal{L}_{b},\end{split}$
where $\overline{p_{\cdot,j}}=\frac{1}{N}\sum_{i=1}^{N}p_{i,j},j\in\{1,...,|\mathcal{Y}|\}$ represents the average predicted probability of the $j$-th target label variable. In this way, we can estimate $H(T)$ via computing the batch entropy $\mathcal{L}_{b}$. It measures the uncertainty or diversity of the predicted probability distribution over the label space. This allows us to introduce a structured regularization term that encourages uniformity across different classes in the latent space. After that, we can balance the learning of different features or labels, preventing the model from overemphasizing certain features or labels that are more prevalent in the training data but may not be representative of the true data distribution.
Structured Probabilistic Coding
We incorporate the structured regularization from the target label space into the probabilistic coding framework, named Structured Probabilistic Coding (SPC). The total objective of SPC can be:
$\begin{split}\mathcal{L}_{SPC}&=\mathcal{L}_{PC}-\gamma H(T)\gtrapprox\mathcal{L}_{PC}-\gamma\mathcal{L}_{b}\\ &=\mathbb{E}_{t\sim p_{\theta}(t|x)}[-\log q(y|t)]+\beta KL(p_{\theta}(t|x);r(t))\\ &\quad+\gamma\sum_{j=1}^{|\mathcal{Y}|}\overline{p_{\cdot,j}}\log\left(\overline{p_{\cdot,j}}\right),\end{split}$
where $\gamma>0$ is a hyperparameter controlling the strength of the regularization. The first two terms combine probabilistic encoding and task prediction into one encoder module with the network $\theta$. The last regularization term promotes class-level uniformity in the latent space under the multivariate Gaussian distribution. Totally, the goal of SPC is to maintain the Gaussian structure of the neighborhood in the input space, as well as achieve the best possible coverage of the hidden space with uniformity across classes.
Applications for Downstream Tasks
We apply the SPC framework to enhance the generalization ability of pre-trained language models for various natural language understanding (NLU) tasks. Due to the ability of learning informative and compact representations, the proposed SPC framework is well-suited for classification and regression tasks. For classification tasks, the lower bound of $I(T;Y)$ can amount to a classic cross-entropy loss (Achille and Soatto 2018; Amjad and Geiger 2019). Similarly, for regression tasks, the lower bound of $I(T;Y)$ can be equivalent to a classic mean squared error loss.
Experiments
Experimental Setups
Dataset Task # Label # Train # Val # Test Classification EmojiEval Emoji prediction 20 45,000 5,000 50,000 EmotionEval Social emotion detection 4 3,257 374 1,421 HatEval Hate speech detection 2 9,000 1,000 2,970 IronyEval Irony detection 2 2,862 955 784 OffensEval Offensive language detection 2 11,916 1,324 860 SentiEval Sentiment analysis 3 45,389 2,000 11,906 StanceEval Stance detection 3 2,620 294 1,249 ISEAR Emotion reaction prediction 7 3,066 767 3,833 MELD Conversational emotion recognition 7 9,989 1,109 2,610 GoEmotions Fine-grained emotion detection 28 36,308 4,548 4,591 Regression STS-B Semantic similarity prediction - 7,000 1,500 1,400 CLAIRE Plausible clarification ranking - 19,975 2,500 2,500
Table 1: The statistics of all datasets.
Methods EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg. ${}^{\dagger}$ 29.30 64.70 36.70 61.70 52.30 62.90 67.30 - - - - ${}^{\dagger}$ 25.80 65.20 50.60 63.10 73.40 62.90 65.40 - - - - ${}^{\dagger}$ 24.70 66.00 52.60 62.80 71.70 58.30 59.40 - - - - GPT-3.5-turbo 6.34$\pm$0.01 73.23$\pm$0.18 48.30$\pm$0.11 66.81$\pm$3.26 63.71$\pm$0.13 40.40$\pm$3.13 39.45$\pm$0.10 67.22$\pm$0.09 41.46$\pm$0.11 25.21$\pm$0.08 47.21 BERT CE 22.30$\pm$0.60 76.05$\pm$1.41 44.67$\pm$1.78 59.38$\pm$3.01 80.16$\pm$1.26 70.54$\pm$0.44 65.21$\pm$0.71 67.17$\pm$0.78 39.80$\pm$0.84 46.29$\pm$0.79 57.16 CE+CP 21.91$\pm$0.71 76.28$\pm$1.20 45.97$\pm$2.93 64.06$\pm$2.41 78.99$\pm$1.57 70.68$\pm$0.31 65.83$\pm$0.39 67.20$\pm$0.95 39.54$\pm$1.61 46.39$\pm$0.63 57.69 CE+AT 22.93$\pm$0.70 75.08$\pm$1.23 46.30$\pm$3.61 64.23$\pm$2.04 79.68$\pm$1.59 70.55$\pm$0.57 66.46$\pm$1.13 65.70$\pm$0.69 39.84$\pm$0.38 47.37$\pm$0.54 57.81 CE+SCL 21.72$\pm$0.51 75.43$\pm$1.37 45.86$\pm$1.15 65.39$\pm$2.46 80.20$\pm$0.56 70.70$\pm$0.79 65.34$\pm$0.60 67.54$\pm$0.64 40.00$\pm$1.96 46.50$\pm$0.46 57.87 VIB 21.31$\pm$0.62 77.37$\pm$0.71 45.99$\pm$1.93 63.82$\pm$1.00 80.37$\pm$1.11 70.39$\pm$0.31 65.43$\pm$0.60 67.24$\pm$0.57 38.52$\pm$0.51 45.89$\pm$1.10 57.63 MINE-IB 21.29$\pm$0.31 76.60$\pm$0.41 47.64$\pm$2.11 65.86$\pm$2.57 78.67$\pm$2.28 69.85$\pm$0.54 65.35$\pm$0.88 67.62$\pm$0.40 41.23$\pm$0.67 46.87$\pm$0.42 58.10 MEIB 21.87$\pm$0.73 76.70$\pm$0.82 48.27$\pm$1.72 65.87$\pm$2.14 80.49$\pm$0.81 70.55$\pm$0.57 65.59$\pm$1.58 67.44$\pm$0.50 39.30$\pm$0.61 46.26$\pm$0.81 58.23 SPC (Ours) 24.19$\pm$1.55 77.15$\pm$0.73 57.48$\pm$2.99 65.85$\pm$1.07 80.65$\pm$0.78 70.74$\pm$0.12 67.17$\pm$1.08 68.94$\pm$0.35 42.68$\pm$0.94 47.62$\pm$1.38 60.25 RoBERTa CE 30.25$\pm$1.32 77.41$\pm$1.33 45.49$\pm$4.70 57.99$\pm$4.96 78.74$\pm$2.20 71.80$\pm$0.93 66.78$\pm$1.34 70.00$\pm$0.45 39.23$\pm$0.41 46.64$\pm$1.15 58.43 CE+CP 31.12$\pm$0.84 77.54$\pm$0.70 48.59$\pm$3.28 58.75$\pm$6.19 79.50$\pm$0.98 72.82$\pm$0.29 66.89$\pm$1.67 70.58$\pm$0.71 40.74$\pm$0.89 47.98$\pm$0.65 59.45 CE+AT 32.00$\pm$0.93 77.30$\pm$1.07 44.71$\pm$4.76 60.17$\pm$3.17 79.81$\pm$1.11 72.51$\pm$0.44 67.81$\pm$0.95 70.97$\pm$0.68 40.10$\pm$0.60 47.89$\pm$1.21 59.33 CE+SCL 31.09$\pm$1.85 76.98$\pm$2.02 49.51$\pm$2.86 60.71$\pm$4.23 80.39$\pm$0.83 73.16$\pm$0.44 66.73$\pm$1.54 70.26$\pm$0.45 40.64$\pm$1.02 47.87$\pm$0.86 59.72 VIB 29.71$\pm$0.79 77.99$\pm$0.86 49.39$\pm$3.08 59.93$\pm$4.57 79.63$\pm$0.66 72.81$\pm$0.39 68.40$\pm$0.52 70.74$\pm$0.44 38.94$\pm$0.55 46.23$\pm$0.18 59.38 MINE-IB 31.70$\pm$0.45 78.79$\pm$0.58 46.39$\pm$2.82 57.39$\pm$8.27 79.76$\pm$0.67 72.85$\pm$0.56 67.27$\pm$1.00 70.15$\pm$0.58 41.80$\pm$2.14 48.88$\pm$1.04 59.50 MEIB 29.94$\pm$1.30 78.73$\pm$0.90 49.34$\pm$2.42 60.54$\pm$2.70 79.68$\pm$0.98 72.78$\pm$0.29 67.89$\pm$1.70 70.86$\pm$0.61 39.00$\pm$0.37 47.18$\pm$1.15 59.59 SPC (Ours) $*{}^{*}$$\pm$0.48 79.01$\pm$0.61 $*{}^{*}$$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 $*{}^{*}$$\pm$0.63 $*{}^{*}$$\pm$0.59 $*{}^{*}$$\pm$0.29 $*{}^{*}$$\pm$0.60 62.47
Table 2: Classification evaluation (%) on 10 benchmark datasets. ${}^{\dagger}$ means the results are from Barbieri et al. (2020). For other methods, we run five random seeds and report the average result on test sets. BERT and RoBERTa are the model backbones for deep representation learning technologies. Best results for each dataset are highlighted in bold. $*{}^{*}$ represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
Methods STS-B CLAIRE Avg. Spearman Pearson Spearman Pearson MSE 88.33$\pm$0.32 88.80$\pm$0.36 47.10$\pm$3.36 48.06$\pm$3.43 68.07 MSE+CP 88.45$\pm$0.43 89.07$\pm$0.32 48.71$\pm$1.71 49.51$\pm$1.59 68.93 VIB 88.45$\pm$0.50 89.01$\pm$0.40 48.11$\pm$1.46 49.07$\pm$1.50 68.66 MEIB 88.61$\pm$0.14 89.13$\pm$0.17 48.97$\pm$1.86 47.87$\pm$1.83 68.64 SPC (Ours) 88.89$\pm$0.30 89.47$\pm$0.31 $*{}^{*}$$\pm$1.54 $*{}^{*}$$\pm$1.58 69.57
Table 3: Regression evaluation (%) on 2 datasets with RoBERTa backbone. For each method, we run five random seeds and report the average result on test sets. $*{}^{*}$ represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
Methods EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg. SPC 32.54$\pm$0.48 79.01$\pm$0.61 59.80$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 69.02$\pm$0.63 71.01$\pm$0.59 43.99$\pm$0.29 50.04$\pm$0.60 62.47   - w/o Structured 30.98$\pm$0.89 78.60$\pm$0.54 56.64$\pm$8.75 62.12$\pm$6.97 79.16$\pm$1.28 72.23$\pm$0.77 68.90$\pm$0.60 70.79$\pm$0.22 43.75$\pm$0.67 46.61$\pm$1.26 60.98   - w/o Structured - w/o Probabilistic 30.25$\pm$1.32 77.41$\pm$1.33 45.49$\pm$4.70 57.99$\pm$4.96 78.74$\pm$2.20 71.80$\pm$0.93 66.78$\pm$1.34 70.00$\pm$0.45 39.23$\pm$0.41 46.64$\pm$1.15 58.43
Table 4: Ablation results (%) on classification tasks. We experiment with RoBERTa backbone.
Datasets and Downstream Tasks
We conduct experiments on various classification and regression tasks, as shown in Table 1. Concretely, following Barbieri et al. (2020), we experiment on 7 classification tasks about tweet analysis on social media, i.e., EmojiEval (Barbieri et al. 2018), EmotionEval (Mohammad et al. 2018), HatEval (Basile et al. 2019), IronyEval (Hee, Lefever, and Hoste 2018), OffensEval (Zampieri et al. 2019), SentiEval (Rosenthal et al. 2014), and StanceEval (Mohammad et al. 2016). To better evaluate the generalization of the method for cross-domain scenes, we also experiment on 3 emotion-related datasets from different domains, i.e., ISEAR (Scherer and Wallbott 1994), MELD (Poria et al. 2019), and GoEmotions (Demszky et al. 2020). Besides, we experiment on 2 regression benchmarks, i.e., STS-B (Cer et al. 2017) and CLAIRE (Roth, Anthonio, and Sauer 2022). See the appendix for more descriptions of datasets and tasks.
Comparison Methods
We compare against the 4 universal models (i.e., SVM, FastText, BiLSTM, and GPT-3.5-turbo) and 7 representative deep representation learning technologies (i.e., CE/MSE, CE/MSE+CP, CE+AT, CE+SCL, VIB, MINE-IB, and MEIB). VIB, MINE-IB, and MEIB belong to probabilistic embedding methods, while the others belong to deterministic embedding methods. For these representation learning technologies, we use pre-trained language models, i.e., BERT (Devlin et al. 2019), and RoBERTa (Liu et al. 2019), as the model backbone for fine-tuning on downstream tasks. Concretely, we use bert-base-uncased [^2] https://huggingface.co/ and roberta-base2 to initialize BERT and RoBERTa for fine-tuning on downstream tasks, respectively.
SVM (Cortes and Vapnik 1995) is a machine learning algorithm with a hinge loss that aims to find the best hyperplane to separate data points into different classes. FastText (Joulin et al. 2017) is an efficient text classification method with negative log-likelihood loss based on n-gram features and a hierarchical softmax. BiLSTM is a bidirectional recurrent neural network (Hochreiter and Schmidhuber 1997) that can be used for classification with cross-entropy loss. GPT-3.5-turbo [^3] https://chat.openai.com is an enhanced generative pre-trained transformer model based on text-davinci-003, optimized for chatting. [^4] We present the zero-shot results of the snapshot from June 13th 2023 based on specific inputs, including task descriptions, task instructions, and evaluation texts.
CE/MSE means a fine-tuned baseline with a cross-entropy (CE) loss for classification tasks or a mean squared error (MSE) loss for regression tasks. CE/MSE+CP (Pereyra et al. 2017) is an entropy regularization method that fits a deterministic network by optimizing an objective that combines the CE/MSE loss with a confidence penalty term. CE+AT (Miyato, Dai, and Goodfellow 2017) uses a cross-entropy objective with classical adversarial training. CE+SCL (Gunel et al. 2021) combines CE and supervised contrastive learning (SCL) (Khosla et al. 2020). SCL allows for multiple positives per anchor, thus adapting contrastive learning to the fully supervised setting. VIB (Alemi et al. 2017; Mahabadi, Belinkov, and Henderson 2021) is an efficient variational estimation method based of the information bottleneck (IB) principle (Tishby and Zaslavsky 2015). MINE-IB (Belghazi et al. 2018) is an neural estimation method of the IB principle with a continuous setting. MEIB (An, Jammalamadaka, and Chong 2023) is a variational approach to stochastic embedding in which maximum conditional entropy acts as the bottleneck. MEIB encourages obvious inputs that can be easily classified to take broader embedding areas by assigning larger entropy.
Evaluation Metrics
We use the same evaluation metric from the original tasks. For the evaluation on classification tasks, the macro-averaged F1 over all classes is applied in most cases. There are three exceptions: stance (macro-averaged of F1 of favor and against classes), irony (F1 of ironic class), and sentiment analysis (macro-averaged recall). Following Barbieri et al. (2020), we report a global metric based on the average of all dataset-specific metrics. For the evaluation on regression tasks, we apply both Pearson and Spearman correlation coefficients. Besides, the $t$-test (Kim 2015) is used to verify the statistical significance of the differences between the results of our SPC and the best non-SPC method on the current dataset.
Implementation Details
All experiments are conducted on a single NVIDIA Tesla A100 80GB card. The validation sets are used to tune hyperparameters and choose the optimal model. For each method, we run five random seeds and report the average result of the test sets. The network parameters are optimized by using Adamax optimizer (Kingma and Ba 2015) with the learning rate of $5e^{-5}$, the weight decay term of $\{0,0.01,0.001\}$. The dropout rate is set to 0.2. The trade-off parameter $\beta$ is searched from $\{0.001,0.01,0.1,1,10\}$, and the parameter $\gamma$ is searched from $\{0.01,0.05,0.1,0.5,1,5,10\}$. Besides, we conduct experiments using an epoch number of $20$, a total batch size of $128$, and a maximum token length of $128$. The maximum patience for early stopping is set to 5 epochs. More experimental details are listed in the Appendix.
Overall Results
Performance on Classification Tasks
The overall results for 10 classification tasks are summarized in Table 2. Our SPC consistently obtains the best average performance over comparison methods. When using BERT and RoBERTa backbones, SPC can enhance average performance by +3.1% and +4.0% compared to CE for all classification tasks, respectively. The results indicate the good generalization ability of our method to unseen test sets and show the superiority of our method on classification tasks. We notice that SPC achieves big improvements for HatEval and IronyEval, i.e., +14.3% macro-F1 scores and +7.3% F1 scores of ironic, respectively. In HateEval, there is a topic distribution disparity between the validation set and the test set. Additionally, the IronyEval task requires complex semantic understanding, with subtle differences between ironic and non-ironic texts. These results indicate that our SPC has a good generalization capability on the above specific scenarios, i.e., topic shifts and subtle semantic labels.
Performance on Regression Tasks
Table 3 presents the overall results of comparison methods in terms of Spearman and Pearson correlation coefficients for two regression tasks. It shows that SPC obtains better regression results on both datasets. Besides, when using RoBERTa backbone, compared to MSE, SPC achieves +1.5% absolute improvements in terms of the average performance. This demonstrates the superiority and generalization of SPC to unseen test sets on regression tasks.
Methods EmotionEval $\rightarrow$ GoEmotions MELD $\rightarrow$ GoEmotions ISEAR $\rightarrow$ GoEmotions Avg. CE 73.79$\pm$2.57 30.71$\pm$0.54 42.99$\pm$2.10 49.16 CE+AT 72.54$\pm$3.89 32.05$\pm$1.69 44.11$\pm$1.44 49.57 VIB 74.73$\pm$3.52 30.50$\pm$1.05 41.88$\pm$1.65 49.03 MEIB 75.55$\pm$2.05 30.11$\pm$1.33 42.10$\pm$0.61 49.25 SPC 77.47$\pm$2.46 33.95$\pm$1.16 44.36$\pm$1.29 51.93
Table 5: Out-of-distribution evaluation results (%). For instance, “EmotionEval $\rightarrow$ GoEmotions” refers to training the model on the training set of EmotionEval and making predictions using the test set of GoEmotions. We experiment with RoBERTa backbone. We run five random seeds and report the average results on test sets of target domains. Labels that do not appear in the training corpus are not evaluated.
Ablation Study
We conduct ablation studies by removing the structured regularization (w/o Structure) and probabilistic coding (w/o Probabilistic). For classification, Table 4 shows the ablation results on all tasks. When removing the structured regularization, SPC w/o Structure obtains inferior performance in terms of all classification metrics. When further removing probabilistic embedding w/o Probabilistic, the results decline significantly. It reveals the effectiveness of structured regularization and probabilistic coding. Since the label space of regression is a one-dimensional real number, our SPC is degraded to probabilistic coding. The ablation w/o Probabilistic MSE is equivalent to the standard MSE. From Table 3, the average performance declines 1.5% on regression metrics, which confirms the effectiveness of probabilistic coding for regression.
Figure 2: Results of different methods against different sizes of training set with RoBERTa backbone.
Generalization Evaluation
We further evaluate the generalization capability of SPC under the following two settings: training with limited data and testing in out-of-distribution (OOD) scenarios.
Comparison under Different Training Size
We experiment under different ratios of the training set to evaluate the generalization when training with limited data. Specifically, given a predefined ratio (e.g., 20%) and a random seed, we randomly sample from the original training set. We obtain 5 training subsets by independently and repeatedly sampling five times from the original training set with 5 different random seeds. All methods are trained on these 5 subsets of the training set, and we report the average results on the test set. Figure 2 shows results of CE, VIB, MEIB, and our SPC against different sizes of training set with RoBERTa backbone. compared to CE, VIB, and MEIB, SPC achieves superior performance on most datasets against different ratios of the training set. It indicates that SPC can enhance the generalization ability of pre-trained language models, even when dealing with limited training data.
Evaluation on Out-of-Distribution
We choose emotion-related benchmarks including EmotionEval, MELD, ISEAR, and GoEmotions, which aim to predict the emotional state but are collected from different domains. To implement OOD scenarios, we train the model on the original training set from a source domain, select the best model based on the validation set of the source domain, and test on the test set of a target domain. To avoid the interference of label mapping bias between different taxonomies, each model is trained on the dataset with coarse-grained taxonomy to predict the label for another dataset with fine-grained taxonomy. Table 5 shows the performance under OOD scenarios. Our SPC obtains the best results on all OOD settings. The fact exhibits SPC’s better generalization capabilities in handling OOD scenarios across different domain shifts. On the one hand, SPC leverages variational inference in the output space, which can better control and utilize randomness and uncertainty of data. On the other hand, SPC introduces the structure information of target task space, making latent space probability distribution better reflect task-related information and generalizing the model to new data.
Methods Noisy EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg. CE 10% 30.66$\pm$0.89 78.15$\pm$0.88 47.06$\pm$5.40 56.90$\pm$4.58 79.46$\pm$0.80 72.36$\pm$0.74 67.39$\pm$1.86 70.40$\pm$0.97 42.01$\pm$1.94 47.85$\pm$1.08 59.22 VIB 10% 30.74$\pm$0.48 77.78$\pm$2.05 47.64$\pm$1.57 58.66$\pm$10.60 79.96$\pm$0.73 72.13$\pm$0.54 67.54$\pm$1.20 70.85$\pm$0.33 38.63$\pm$0.89 47.30$\pm$1.65 59.12 MEIB 10% 31.02$\pm$0.47 78.94$\pm$0.46 49.28$\pm$4.58 57.21$\pm$8.07 80.19$\pm$0.83 72.09$\pm$0.68 68.26$\pm$0.68 70.85$\pm$0.38 38.67$\pm$0.97 46.93$\pm$1.06 59.34 SPC 10% 32.25$\pm$0.69 78.88$\pm$0.47 56.13$\pm$5.36 58.88$\pm$4.94 80.14$\pm$0.28 72.76$\pm$0.06 68.57$\pm$1.01 71.10$\pm$0.62 43.90$\pm$1.13 49.32$\pm$1.22 61.19 CE 20% 31.96$\pm$0.88 77.01$\pm$1.51 49.12$\pm$0.72 60.82$\pm$3.56 79.54$\pm$1.64 72.06$\pm$0.63 68.49$\pm$1.20 70.32$\pm$0.26 40.16$\pm$1.94 47.78$\pm$0.84 59.73 VIB 20% 30.46$\pm$0.59 79.00$\pm$0.49 47.91$\pm$2.20 60.67$\pm$4.82 79.15$\pm$1.22 72.26$\pm$0.29 66.83$\pm$0.52 71.02$\pm$0.25 39.33$\pm$1.47 47.83$\pm$1.38 59.45 MEIB 20% 30.84$\pm$0.75 78.38$\pm$0.88 50.02$\pm$5.18 55.12$\pm$7.07 78.17$\pm$2.55 71.63$\pm$1.11 68.05$\pm$0.81 70.68$\pm$0.38 39.09$\pm$0.87 47.29$\pm$1.22 58.93 SPC 20% 32.51$\pm$0.83 77.97$\pm$1.12 55.41$\pm$6.00 66.40$\pm$4.26 80.33$\pm$0.48 72.50$\pm$0.55 68.89$\pm$1.60 71.10$\pm$0.39 43.96$\pm$0.50 49.04$\pm$0.42 61.81 CE 30% 31.82$\pm$0.75 77.61$\pm$0.90 50.69$\pm$2.80 58.90$\pm$11.45 78.11$\pm$2.07 70.15$\pm$0.5 69.07$\pm$1.07 70.74$\pm$0.56 40.61$\pm$2.06 47.76$\pm$2.29 59.55 VIB 30% 30.85$\pm$0.53 78.23$\pm$0.79 48.22$\pm$1.97 58.81$\pm$8.84 79.38$\pm$0.62 72.15$\pm$0.52 67.59$\pm$0.93 70.27$\pm$0.74 38.71$\pm$1.19 47.16$\pm$1.32 59.14 MEIB 30% 30.74$\pm$0.87 77.99$\pm$0.69 49.98$\pm$4.00 57.57$\pm$5.19 72.53$\pm$5.53 71.83$\pm$0.40 67.88$\pm$0.68 69.86$\pm$1.24 39.39$\pm$1.06 47.43$\pm$1.52 58.52 SPC 30% 32.27$\pm$0.48 78.13$\pm$1.13 56.04$\pm$7.44 59.27$\pm$8.56 80.32$\pm$0.53 72.44$\pm$0.36 69.77$\pm$0.93 70.91$\pm$0.30 43.29$\pm$0.53 49.82$\pm$2.55 61.23
Table 6: Results (%) against different ratios of label noises. RoBERTa is applied as the model backbone.
Figure 3: Clustering performances of the output representations learned by different optimization objectives. Silhouette coefficient (SC) and adjusted rand index (ARI) are used to measure data-related and task-related clustering abilities, respectively. We experiment with RoBERTa backbone.
Robustness Evaluation
We experiment to demonstrate the robustness by assessing how well the models can handle noisy labels. It is crucial for real-world scenarios where data can often be unreliable. Specifically, we randomly choose 10%, 20%, 30% of training data and flip their labels to any category randomly with equal probability. We run experiments five times and compute the mean and standard variance of the results. As shown in Table 6, under all settings, SPC consistently outperforms CE, VIB, and MEIB. It indicates that SPC performs more robustly on noisy training data. Besides, compared to CE, SPC improves average performance by +2.0%, +2.1%, and +1.7% with noise ratio of 10%, 20%, and 30% on classification tasks. The results prove that SPC can better control and utilize randomness and uncertainty of data.
Representation Quality Evaluation
To assess the quality of the representations, we evaluate the clustering performance of output representations obtained by different optimization objectives. Following Hu et al. (2023), we apply silhouette coefficient (SC) and adjusted rand index (ARI) to measure the clustering ability relevant to input data and target labels, respectively. Figure 3 shows SC and ARI of representations learned by various learning objectives. According to the results, SPC achieves higher ARI or SC values compared to other objectives (CE, VIB, and MEIB) across most datasets. It suggests that SPC effectively achieves a balance between data compression and task encoding, thereby promoting the generalization of pre-trained language models for downstream tasks.
Related Work
According to the nature of embeddings, representation learning approaches can be broadly categorized into deterministic embedding and probabilistic embedding.
Deterministic Embedding
Deterministic embedding maps each data point to a fixed vector. The representative works include entropy regularization (Pereyra et al. 2017), adversarial training (Miyato, Dai, and Goodfellow 2017), and contrastive learning (Gunel et al. 2021; Hu et al. 2023).
Probabilistic Embedding
Probabilistic embedding (Vilnis and McCallum 2015) learns a probability distribution that maps each data point to a distribution. The probabilistic approach can better capture the complexity and uncertainty of data, handle redundant information, and provide better discriminative representations. It has been applied to various domains such as computer vision (Oh et al. 2019; Shi and Jain 2019; Chang et al. 2020; Ramé and Cord 2021) and natural language processing (Mahabadi, Belinkov, and Henderson 2021; Wang et al. 2021; Hu et al. 2022a).
Most probabilistic embedding methods (Kingma and Welling 2014; Alemi et al. 2017; Higgins et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) are or can be built upon the principle of information bottleneck (IB) theory (Tishby, Pereira, and Bialek 1999; Tishby and Zaslavsky 2015). The principle aims to find a maximally compressed representation of the input that maximally preserves information about the output, striking a balance between compression and prediction. VIB (Alemi et al. 2017) is an efficient variational estimation method of the IB principle. For tractable application of IB in a continuous setting, Belghazi et al. (2018) propose a mutual information neural estimation method with IB principle, denoted as MINE-IB. And Ragonesi et al. (2021) employ MINE-IB to learn unbiased representations. Fischer (2020) and Ramé and Cord (2021) introduce a conditional mutual information term to alleviate the over- or under-compression issue of traditional IBs. Moreover, variational autoencoder (VAE) (Kingma and Welling 2014) is a special case of an unsupervised VIB and can be used to encourage disentangled representations (Higgins et al. 2017). Hu et al. (2022a) apply VAE to the masked pre-training process for learning diverse and well-formed contextual representations. Recently, An, Jammalamadaka, and Chong (2023) use the conditional entropy of the stochastic embedding as a confidence indicator and encourage the model to assign larger variance to more certain inputs.
Conclusion
This paper proposes a new Structured Probabilistic Coding (SPC) framework to enhance the generalization ability of pre-trained language models for better language understanding. SPC adopts an encoder-only probabilistic coding technology to simultaneously perform information compression and task prediction. And a structured regularization is introduced to control probability distribution and promote class-level uniformity in the latent space. Experiments on 12 benchmarks show that SPC achieves the best performance on various classification and regression tasks. The results demonstrate that SPC can enhance pre-trained language models’ generalization capability, robustness to label noise, and the clustering quality of output representations.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (No. 2022YFC3302102). The authors thank the anonymous reviewers and the meta-reviewer for their helpful comments on the paper.
Appendix Overview
In this supplementary material, we provide: (i) a detailed description of experimental setups, and (ii) supplementary experiments.
Appendix A Experimental Setups
Datasets and Downstream Tasks
We conduct extensive experiments on various natural language understanding tasks including 10 classification tasks, and 2 regression tasks. The descriptions of each dataset and task are listed as follows:
Classification Tasks
  • EmojiEval (Barbieri et al. 2018) is designed for emoji prediction, which aims to predict its most likely emoji given a tweet. Its label set comprises 20 different emoji.
  • EmotionEval (Mohammad et al. 2018) involves predicting the most probable emoji of a tweet and is based on Emoji Prediction challenge conducted during SemEval-2018.
  • HatEval (Basile et al. 2019) stems from SemEval-2019 Hateval challenge and is used to predict whether a tweet is hateful towards immigrants or women.
  • IronyEval (Hee, Lefever, and Hoste 2018) is from SemEval-2018 Irony Detection and consists of identifying whether a tweet includes ironic intents or not.
  • OffensEval (Zampieri et al. 2019) is from SemEval-2019 OffensEval and involves predicting if a tweet contains any form of offensive language.
  • SentiEval (Rosenthal et al. 2014) comes from SemEval 2017 and includes data from previous runs (2013, 2014, 2015, and 2016) of the same SemEval task. The goal is to determine if a tweet is positive, negative, or neutral.
  • StanceEval (Mohammad et al. 2016) involves determining if the author of a piece of text has a favorable, neutral, or negative position towards a proposition or target.
  • ISEAR (Scherer and Wallbott 1994) is from International Survey On Emotion Antecedents And Reactions project and contains reports on seven emotions each by close to 3000 respondents in 37 countries on all 5 continents. It aims to predict the emotion reaction.
  • MELD (Poria et al. 2019) contains multi-party conversation videos collected from Friends TV series, where two or more speakers are involved in a conversation. It is used to detect emotions in each utterance. [^5] The MELD dataset contains many types of context, including dialogue, speaker, and multi-modal data. Different from other task-oriented methods (Hu, Wei, and Huai 2021; Hu et al. 2022b), this work only considers the context-free textual utterance to better evaluate sentence classification performance.
  • GoEmotions (Demszky et al. 2020) is a corpus of comments from Reddit, with human annotations to 27 emotion categories or neutral. It is used for fine-grained emotion detection. [^6] The GoEmotions dataset contains nearly 16% multi-label data. In this paper, we remove this portion to better evaluate the multi-class classification performance.
Hyperparameter EmojiEval EmotionEval HatEval IronyEval OffensEval Trade-off weight $\beta$ 0.1 0.1 10 1 1 Trade-off weight $\gamma$ 10 0.1 0.1 10 0.1 Weight decay 0 0 0.001 0 0 Normalization False True False True False Hyperparameter SentiEval StanceEval ISEAR MELD GoEmotions Trade-off weight $\beta$ 0.01 1 0.1 1 0.001 Trade-off weight $\gamma$ 10 0.5 0.1 0.05 0.01 Weight decay 0 0 0.001 0 0 Normalization False False False True False
Table 7: Hyperparameters of the proposed SPC with RoBERTa backbone on classification tasks.
Hyperparameter STS-B CLAIRE Trade-off weight $\beta$ 0.1 0.01 Weight decay 0 0 Normalization False False
Table 8: Hyperparameters of the proposed SPC with RoBERTa backbone on regression tasks.
Figure 4: Performance against different trade-off weights $\beta$ of probabilistic coding for classification tasks. The experiments are conducted based on RoBERTa backbone. The grey line represents the results of CE baseline.
Figure 5: Performance of the optimal trade-off weight $\gamma$ for classification tasks. We experiment with RoBERTa backbone. The Y-axis refers to relative improvements between SPC and its variant SPC w/o Structure.
Regression Tasks
  • STS-B (Cer et al. 2017) is a collection of sentence pairs drawn from news headlines, video and image captions, and natural language inference data. The semantic similarity task is to predict the semantic similarity score from 1 (very dissimilar) to 5 (very similar) given each pair.
  • CLAIRE (Roth, Anthonio, and Sauer 2022) datasets consists of manually clarified how-to guides from wikiHow [^7] https://www.wikihow.com/ with generated alternative clarifications and human plausibility judgements. The goal of plausible clarifications ranking task is to predict the continuous plausibility score on a scale from 1 (very implausible) to 5 (very plausible) given the clarification and its context.
Implementation Details
We report the detailed hyperparameter settings of SPC with RoBERTa backbone in Table 7 and Table 8. In the implementation of SPC, the hidden vector represents the output representation, and its dimension size is consistent with the dimension size of label space in each task. For each baseline, we fine-tune the key parameters following the original paper for fair comparison and to obtain corresponding optimal performance.
Appendix B Supplementary Experiments
Parameter Analysis
In this part, we analyze what the trade-off parameters $\beta$ and $\gamma$ control in our SPC.
Performance against different weights of probabilistic coding
Figure 4 shows results against different values of $\beta$. With the enhancement of the optimization strength of probabilistic coding ($\beta\uparrow$), SPC is prone to assign larger variance for noisy samples and small variance for high quality ones.
Performance against optimal weight of task expression
Figure 5 shows relative improvements between SPC and its ablated variant (i.e., w/o Structured) against the optimal $\gamma$. By introducing task-related underlying structured patterns in representation learning ($\gamma>0$), SPC achieves varying degrees of relative improvements on all tasks, particularly in HatEval and IronyEval. A larger value of $\gamma$ indicates that this type of task requires enhanced task-related learning ability.